### 24 Mar 2026 v1

# Hook-decomposable modules and their resolutions

Abstract
We compare several classes of biparameter persistence modules: $\gamma$-products of monoparameter modules, hook-decomposable modules, modules admitting a Smith-type structure theorem, and modules of projective dimension 1. We determine all logical implications among these classes, providing explicit counterexamples showing that the converses fail when appropriate. In particular, $\gamma$-products (i.e., hook-decomposable modules) form a very small subclass of biparameter modules, precisely the ones for which a structure theorem still holds, thus making explicit the richer structural complexity of the biparameter setting compared to the monoparameter one.
1 Motivation and contribution
Persistent homology is a fundamental tool in topological data analysis, with monoparameter persistence extensively studied and applied [8, 4, 6, 7]. Multiparameter persistence, while offering greater discriminative power [5, 9], remains less commonly used, partly due to the lack of a simple structural interpretation of its capabilities. A natural way to investigate this difference could be to examine biparameter modules obtained from constructions on monoparameter ones, in order to understand how such “simple” biparameter modules sit inside the broader category.
In previous work [10], we introduced one such construction, the $\gamma$-product, and showed that the resulting biparameter modules coincide precisely with hook-decomposable modules. Motivated by this, in the present work we analyze the logical relations between four classes of biparameter persistence modules naturally arising from, or closely related to, this construction: $(i)$ products of monoparameter modules (i.e., $\gamma$-products), $(ii)$ hook-decomposable modules, $(iii)$ modules for which a Smith-type structure theorem applies, $(iv)$ modules of projective dimension 1, $(v)$ modules of projective dimension 1 whose minimal free resolution splits. We show that $(i)$, $(ii)$, and $(iii)$ are equivalent, and that $(v)$ implies $(ii)$, which in turn implies $(iv)$; none of these last two implications has a converse. The resulting diagram of implications is given in Figure 1, while a schematic arrangement of these classes, as it will follow from the implications proved below, is shown in Figure 2.
Figure 1: Implication diagram that we establish in this work.
Figure 2: Visual representation of the studied classes. Examples of possible supports are shown for each one. Lines: green = births; red-dotted = deaths; orange-dashed = merges. Blue areas = supports interior.
2 Background notions
Let $\mathbb{F}$ be a field and $M$ be a biparameter persistence module, i.e., a finitely generated, standard bigraded $\mathbb{F}[x,y]$-module. We recall the notions relevant to our analysis.
Definition 2.1.
Let $M$ be as above.
  1. 1.
    $\gamma$-product. A construction producing a biparameter module from two monoparameter modules. (We refer to [10] for the precise definition.)
  2. 2.
    Hook-decomposable modules. We call $M$ hook-decomposable if it is a direct sum of hook submodules, i.e., interval modules II for which there exist bigrades $\mathbf{p}\leq\mathbf{q}$, with $\mathbf{q}$ possibly having infinite coordinates, such that its support [^1] The support of a (multi)graded module is the set of (multi)degrees in which the module is nonzero. is of the form $Supp(I)=\{\boldsymbol{\alpha}\in\mathbb{N}^{2}\thinspace|\thinspace\mathbf{p}\leq\boldsymbol{\alpha},\boldsymbol{\alpha}\not\geq\mathbf{q}\}.$ Such class of modules is relevant due to the fact that the rank exact resolution of any module always consists of hook-decomposable modules [2, 1, 3].
  3. 3.
    Structure theorem. We say that $M$ satisfies the structure theorem if it decomposes as
    $M\;\cong\;\bigoplus_{k=1}^{f}\mathbb{F}[x,y](-\mathbf{p}_{k})\;\oplus\;\bigoplus_{j=1}^{t}\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}_{j}-\mathbf{p}_{j}})\big)(-\mathbf{p}_{j}),$
    for some bigrades $\mathbf{p}_{k},\mathbf{p}_{j},\mathbf{q}_{j}$, where $\mathbf{x}^{\,\mathbf{q}_{j}-\mathbf{p}_{j}}=x^{q_{jx}-p_{jx}}y^{q_{jy}-p_{jy}}$, with $\mathbf{p}_{j}=(p_{jx},p_{jy}),\mathbf{q}_{j}=(q_{jx},q_{jy})$.
    Equivalently, $M$ admits a Smith diagonal form.
  4. 4.
    Resolution. A (minimal) graded free resolution of $M$ has the form
    $0\longrightarrow F^{\beta_{2}}\longrightarrow F^{\beta_{1}}\longrightarrow F^{\beta_{0}}\longrightarrow M\longrightarrow 0,$
    where each free module is $F^{\beta_{i}}\;:=\;\bigoplus_{j}\mathbb{F}[x,y]^{\beta_{i,j}}(-\mathbf{p}_{i,j}),$ the integers $\beta_{i,j}$ are the graded Betti numbers of $M$, and $\beta_{i}\;:=\;\sum_{j}\beta_{i,j}.$ Each $\mathbf{p}_{i,j}$ denotes the bidegree of the $j$-th generator in $F^{\beta_{i}}$.
  5. 5.
    Projective dimension. The module $M$ has projective dimension 1 if and only if $\beta_{2}=0$ in its minimal graded free resolution.
  6. 6.
    Split exact sequence. A short exact sequence $0\to A\to B\to C\to 0$ splits if $B\cong A\oplus C$, equivalently if the injection admits a retraction or the surjection admits a section.
3 Hook-decomposability: sufficient and necessary conditions
Theorem 3.1.
(Characterization of hook-decomposability) The following are equivalent: (i)M(i)\,M is a $\gamma$-product; (ii)M(ii)\,M is hook-decomposable; $(iii)$ the structure theorem holds for $M$.
Sketch of the proof.
The equivalence between $(i)$ and $(ii)$ was established in [10].
For the equivalence between $(ii)$ and $(iii)$, observe that a hook summand is by definition of the form $\mathbb{F}[x,y](-\mathbf{p})$ or $\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})$, where $\mathbf{p}\leq\mathbf{q}$ are the bigrades determining its support as in Definition 2.1 ($\mathbf{q}$ has infinite coordinates in the first case). Thus, $M$ is hook-decomposable if and only if it decomposes as in (3), that is, if and only if the structure theorem holds for $M$. ∎
Theorem 3.2.
(Properties of the resolutions of hook-decomposable modules)
  1. 1.
    If $M$ has projective dimension 1 and split resolution, then $M$ is hook-decomposable.
  2. 2.
    If $M$ is hook-decomposable, then $M$ has projective dimension 11.
Sketch of the proof.
  1. 1.
    From the hypothesis it follows that $M$ is composed by null or free summands, hence hook summands.
  2. 2.
    From Theorem 3.1, it follows that each hook summand is of the form $\mathbb{F}[x,y](-\mathbf{p})$ or $\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})$ and then their resolutions are $0\rightarrow\mathbb{F}[x,y](-\mathbf{p})\rightarrow\mathbb{F}[x,y](-\mathbf{p})\rightarrow 0$ and $0\rightarrow\mathbb{F}[x,y](-\mathbf{q})\xrightarrow{\;\mathbf{x}^{\,\mathbf{q}-\mathbf{p}}\;}\mathbb{F}[x,y](-\mathbf{p})\rightarrow\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})\rightarrow 0,$ respectively.
Remark $3.$
We now exhibit two examples, included in Figure 2, that show that the converse implications do not hold.
  1. 1.
    The module with one generator $g$ in bidegree $(0,0)$ and the relation xygxyg in bidegree $(1,1)$ is hook-decomposable and has projective dimension 1, yet its resolution does not split , since the injection admits no linear retraction.
  2. 2.
    The module with generators $g$ in bidegree $(0,1)$ and $h$ in bidegree $(1,0)$ and the relation xgyhxg-yh in bidegree $(1,1)$ has projective dimension 1, yet it is not hook-decomposable.
Acknowledgements
This work was carried out within the framework of the following projects: the Programma Regionale Fondo Sociale Europeo+ 2021–2027, Priorità 2 – Istruzione e Formazione – ESO 4.6 (OS-f), with the contribution of Rulex Innovation Labs S.r.l.; the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.5, by the European Union – NextGenerationEU and the Ministry of University and Research (MUR); the National Center for HPC, Big Data and Quantum Computing (CN0000013); Italian National Biodiversity Future Center (NBFC) - National Recovery and Resilience Plan (NRRP) funded by the European Union - NextGenerationEU (project code CN 00000033); Innovation Ecosystem Robotics and AI for Socio-economic Empowerment (RAISE)National Recovery and Resilience Plan (NRRP) funded by the European Union – NextGenerationEU (project code ECS 00000035).

### 15 Apr 2026 v2

@@
 Abstract
-We compare several classes of biparameter persistence modules: $\gamma$-products of monoparameter modules, hook-decomposable modules, modules admitting a Smith-type structure theorem, and modules of projective dimension 1.
+We compare several classes of biparameter persistence modules: $\gamma$-products of monoparameter modules, hook-decomposable modules, modules admitting a Smith-type structure theorem, and modules of projective dimension at most 1.
 We determine all logical implications among these classes, providing explicit counterexamples showing that the converses fail when appropriate.
@@
 In previous work [10], we introduced one such construction, the $\gamma$-product, and showed that the resulting biparameter modules coincide precisely with hook-decomposable modules.
-Motivated by this, in the present work we analyze the logical relations between four classes of biparameter persistence modules naturally arising from, or closely related to, this construction: $(i)$ products of monoparameter modules (i.e., $\gamma$-products), $(ii)$ hook-decomposable modules, $(iii)$ modules for which a Smith-type structure theorem applies, $(iv)$ modules of projective dimension 1, $(v)$ modules of projective dimension 1 whose minimal free resolution splits.
+Motivated by this, in the present work we analyze the logical relations between four classes of biparameter persistence modules naturally arising from, or closely related to, this construction: $(i)$ products of monoparameter modules (i.e., $\gamma$-products), $(ii)$ hook-decomposable modules, $(iii)$ modules for which a Smith-type structure theorem applies, $(iv)$ modules of projective dimension 1, $(v)$ free modules.
 We show that $(i)$, $(ii)$, and $(iii)$ are equivalent, and that $(v)$ implies $(ii)$, which in turn implies $(iv)$; none of these last two implications has a converse.
@@
   5. 5.
-    Projective dimension. The module $M$ has projective dimension 1 if and only if $\beta_{2}=0$ in its minimal graded free resolution.
-  6. 6.
-    Split exact sequence. A short exact sequence $0\to A\to B\to C\to 0$ splits if $B\cong A\oplus C$, equivalently if the injection admits a retraction or the surjection admits a section.
+    Projective dimension. The module $M$ has projective dimension $\leq 1$ if and only if $\beta_{2}=0$ in its minimal free resolution.
+Remark $2.$
+Please note that freeness (condition $(v)$) is equivalent to projective dimension $\leq 0$ and to the minimal free graded resolution being split short exact.
 3 Hook-decomposability: sufficient and necessary conditions
@@
   1. 1.
-    If $M$ has projective dimension 1 and split resolution, then $M$ is hook-decomposable.
+    $(v)\Rightarrow(ii)$. If $M$ is free, then $M$ is hook-decomposable.
   2. 2.
-    If $M$ is hook-decomposable, then $M$ has projective dimension 11.
+    $(ii)\Rightarrow(iv)$. If $M$ is hook-decomposable, then $M$ has projective dimension $\leq 1$.
 Sketch of the proof.
+  1. Each free summand of $M$ is either null or has a relation at infinity.
   2. 1.
-    From the hypothesis it follows that $M$ is composed by null or free summands, hence hook summands.
-  2. 2.
-    From Theorem 3.1, it follows that each hook summand is of the form $\mathbb{F}[x,y](-\mathbf{p})$ or $\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})$ and then their resolutions are $0\rightarrow\mathbb{F}[x,y](-\mathbf{p})\rightarrow\mathbb{F}[x,y](-\mathbf{p})\rightarrow 0$ and $0\rightarrow\mathbb{F}[x,y](-\mathbf{q})\xrightarrow{\;\mathbf{x}^{\,\mathbf{q}-\mathbf{p}}\;}\mathbb{F}[x,y](-\mathbf{p})\rightarrow\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})\rightarrow 0,$ respectively.
+    From Theorem 3.1, it follows that each hook summand is of the form $\mathbb{F}[x,y](-\mathbf{p})$ or
+    $\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})$, so their resolutions are $0\rightarrow\mathbb{F}[x,y](-\mathbf{p})\rightarrow\mathbb{F}[x,y](-\mathbf{p})\rightarrow 0$ and $0\rightarrow\mathbb{F}[x,y](-\mathbf{q})\xrightarrow{\;\mathbf{x}^{\,\mathbf{q}-\mathbf{p}}\;}\mathbb{F}[x,y](-\mathbf{p})\rightarrow\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})\rightarrow 0,$ respectively.
 Remark $3.$
@@
   1. 1.
-    The module with one generator $g$ in bidegree $(0,0)$ and the relation xygxyg in bidegree $(1,1)$ is hook-decomposable and has projective dimension 1, yet its resolution does not split , since the injection admits no linear retraction.
+    The module with one generator $g$ in bidegree $(0,0)$ and the relation xygxyg in bidegree $(1,1)$ is hook-decomposable yet its not free.
+Notably, its resolution does not split, as the injection admits no graded retraction.
   2. 2.
@@
 This work was carried out within the framework of the following projects: the Programma Regionale Fondo Sociale Europeo+ 2021–2027, Priorità 2 – Istruzione e Formazione – ESO 4.6 (OS-f), with the contribution of Rulex Innovation Labs S.r.l.; the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.5, by the European Union – NextGenerationEU and the Ministry of University and Research (MUR); the National Center for HPC, Big Data and Quantum Computing (CN0000013); Italian National Biodiversity Future Center (NBFC) - National Recovery and Resilience Plan (NRRP) funded by the European Union - NextGenerationEU (project code CN 00000033); Innovation Ecosystem Robotics and AI for Socio-economic Empowerment (RAISE)National Recovery and Resilience Plan (NRRP) funded by the European Union – NextGenerationEU (project code ECS 00000035).
+free modules